OpenStudy (anonymous):

determine the resulting ionic formula and names for the following pairs of cations/anions: a.) Chromium (III)/ Sulfate b.) Calcium/Nitrogen

OpenStudy (paxpolaris):

\[\_Cr^{3+} + \_\left( SO_4 \right)^{2-}\] balance the charges..

OpenStudy (paxpolaris):

\[2Cr^{3+} + 3\left( SO_4 \right)^{2-} \to Cr_2\left( SO_4 \right)_3\]

OpenStudy (paxpolaris):

\[3Ca^{2+}+2N^{3-} \to Ca_3N_2\]